OpenStudy (anonymous):

-7s-5=-6s solve for s PLEASE

OpenStudy (anonymous):

i need somee help guys!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (diyadiya):

Ok Add 6s on both sides can you?

OpenStudy (diyadiya):

@barbiejazzybabii can you do that??????

OpenStudy (anonymous):

yes @Diyadiya

OpenStudy (diyadiya):

ok what do you get now?

OpenStudy (anonymous):

i got -1s-5=s??? @Diyadiya

OpenStudy (diyadiya):

Nope -7s-5=-6s -7s - 5 + 6s = -6s + 6s -s -5 = 0 is this clear?

OpenStudy (anonymous):

yes

OpenStudy (diyadiya):

Now can you add 5 on both sides?

OpenStudy (anonymous):

now add 5?

OpenStudy (diyadiya):

Yes!! :D

OpenStudy (anonymous):

now is the answer 5?

OpenStudy (diyadiya):

how>?

OpenStudy (anonymous):

because 0 + 5 = 5

OpenStudy (diyadiya):

what happened to the other side? ;)

OpenStudy (anonymous):

its eliminated?

OpenStudy (diyadiya):

-s -5 = 0 -s -5 + 5 = 0+5 -s = 5 you have to find s not -s :) so what do you do now?

OpenStudy (anonymous):

nothing............isnt it 5?

OpenStudy (diyadiya):

no -s & s are not the same to get s you have to multiply -1 on both sides of -s = 5

OpenStudy (anonymous):

so its -5?

OpenStudy (diyadiya):

Yes! Right :)

OpenStudy (anonymous):

yayy

OpenStudy (diyadiya):

^_^